OpenStudy (anonymous):

As circle has a radius of 3 cm. what happens to the area if the radius is doubled to 6 cm? A.it is doubled B. it is tripled C.it is quadrupled D. it says the same

OpenStudy (anonymous):

a becaouse in a circle if you double one they all double

OpenStudy (anonymous):

k i'm going to see some other answers about this question k

OpenStudy (anonymous):

okay =) that will be good

OpenStudy (anonymous):

to see what other people thing

OpenStudy (anonymous):

No. Area is proportional to the square of the radius.

OpenStudy (anonymous):

So what dows that mean @douglaswinslowcooper ?

OpenStudy (anonymous):

i ment does

OpenStudy (anonymous):

A2/A1 = (r2/r1)^2 A and r are the areas and radii

OpenStudy (anonymous):

Is it because there the variables?? @douglaswinslowcooper

OpenStudy (lyrae):

\[A = \pi r ^{2}\] \[Area1 = \pi \times 3^{2} = 9 \pi\]\[Area2 = \pi \times 6^{2} = 36 \pi\] \[\frac{ area2 }{ area1 } = \frac{ 36 pi }{ 9 pi } = 4\] Conclusion: the area is four times grater in this case.

OpenStudy (anonymous):

Yes Area = pi (radius)^2 New area = pi (2 old radius)^2 = pi 4 (old radius )^2 = 4 old area When something is squared, doubling it quadruples the square.

OpenStudy (anonymous):

So Im guessing @Lyrae and @douglaswinslowcooper That its C.

OpenStudy (anonymous):

Yes. Not a guess.

OpenStudy (lyrae):

Yes.

OpenStudy (anonymous):

Thankyou for taking me step by step in understanding it...and thanks for your time:))) have a nice day:))